If 120.0 g of CH4 is mixed with 120.0 g of H2O, how many grams of H2 is produced?
1
Expert's answer
2020-05-11T14:23:48-0400

"CH_4+H_2O\to CO+3H_2"

As per the above equation,

"1" mole of methane reacts with "1" mole of water to produce "3" moles of Hydrogen.

Moles of methane in "120\ g=\frac{120}{16}=7.5"

Moles of water in "120\ g=\frac{120}{18}=6.66"

So, water is a limiting reagent and hydrogen formation will depend on mass of water reacted.

Moles of hydrogen reacted"=3\times \frac{120}{18}=20"

Mass of "H_2" produced"=2\times 20= 40 \ g"